# Nilpotent structures of oriented neutral vector bundles

### v2

@@ 4 Special nilpotent structures
Suppose that $\nabla$ is flat. Let $K$ be a Lie subgroup of $G$. Then for a $K$-nilpotent structure of $(E,h,\nabla)$, there exists a $K$-admissible frame field $e$ of $N$ on a neighborhood of each point of $M$ such that each local section of $E$ which appears in $e$ is parallel with respect to $\nabla$.
@@
Let $W$ be a $2n$-dimensional subspace of $E^{4n}_{2n}$ spanned by $\eta_{1},\dots,\eta_{n},\eta_{n+1},\dots,\eta_{2n}$. Then $W$ is a light-like subspace of $E^{4n}_{2n}$. Let $SO(2n,2n)_{W}$ be a Lie subgroup of $SO(2n,2n)$ which consists of elements preserving $W$. Then we have $G\subset SO(2n,2n)_{W}$ and $G\not=SO(2n,2n)_{W}$ ([3]). A $SO(2n,2n)_{W}$-nilpotent structure of $(E,h,\nabla)$ is just an $\varepsilon$-nilpotent structure of $(E,h)$ satisfying the Walker condition. If $K=SO(2n,2n)_{W}$, then the conclusion of Proposition 4.2 is not necessarily valid, because of Example 3.1 $\sim$ Example 3.4, while for an almost $\varepsilon$-nilpotent structure $N$ of $E^{4n}_{2n}$ satisfying the Walker condition with respect to the Levi-Civita connection, $\pi_{N}$ gives a constant $2n$-dimensional light-like subspace of $E^{4n}_{2n}$.
@@
See [5], [14] for paraquaternionic structures (notice the signature in (i)). We say that $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$ if $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ satisfy (i), (ii) and
@@
    a section of $E_{I,J_{1},J_{2}}$ given by $N_{r,\theta}:=r(I-(\sin\theta)J_{1}+(\cos\theta)J_{2})$ for $r\in\mbox{\boldmath{$R$}}\setminus\{0\}$ and $\theta\in[0,2\pi)$ is an $H$-nilpotent structure of $(E,h);$
@@
    There exists a unique light-like subbundle $\pi^{*}_{N}$ of $(E,h)$ of rank $2n$ which is locally spanned by
@@
    for any $H$-admissible frame field $e$ of $N$. In particular, $E=\pi_{N}\oplus\pi^{*}_{N}$.
@@
    There exists a unique $\varepsilon$-nilpotent and $H$-nilpotent structure $N^{*}$ of $(E,h)$ such that each $H$-admissible frame field of $N^{*}$ is given by $e^{\prime}=(e_{1},\dots,e_{2n},-e_{2n+1},\dots,-e_{4n})$ for an $H$-admissible frame field $e$ of $N$. In particular, $\pi_{N^{*}}=\pi^{*}_{N}$.
@@
    If $N$ is an $H$-nilpotent structure of $(E,h,\nabla)$, then $N^{*}$ is also an $H$-nilpotent structure of $(E,h,\nabla)$.
Proof  For an $H$-admissible frame field $e$ of $N$, the subbundle of the restriction of $E$ on the domain of $e$ spanned by $\xi^{\prime}_{i}$, $\xi^{\prime}_{n+i}$ ($i=1,\dots,n$) as in (4.4) does not depend on the choice of $e$, because of the definition of $H$ in (4.1). Therefore there exists a unique light-like subbundle $\pi^{*}_{N}$ of $(E,h)$ of rank $2n$ satisfying (a) in Theorem 4.5. Hence we obtain (a) in Theorem 4.5. For $N$, we can find an open cover $\{U_{\lambda}\}_{\lambda\in\Lambda}$ of $M$ and a family $\{e_{\lambda}\}_{\lambda\in\Lambda}$ of $H$-admissible frame fields of $N$ satisfying (i), (ii) in Definition 4.1 for $K=H$. We set
@@
for $\lambda$, $\mu\in\Lambda$ satisfying $U_{\lambda}\cap U_{\mu}\not=\emptyset$. Then $A^{\prime}_{\lambda\mu}$ satisfies $e^{\prime}_{\mu}I^{\prime}_{4n,\varepsilon}=e^{\prime}_{\lambda}I^{\prime}_{4n,\varepsilon}A^{\prime}_{\lambda\mu}$ on $U_{\lambda}\cap U_{\mu}$, and since $A_{\lambda\mu}$ is valued in $H$, $A^{\prime}_{\lambda\mu}$ is also valued in $H$. Therefore there exists an $\varepsilon$-nilpotent and $H$-nilpotent structure $N^{*}$ of $(E,h)$ such that $e^{\prime}_{\lambda}$ ($\lambda\in\Lambda$) are $H$-admissible frame fields of $N^{*}$ and then $N^{*}$ is uniquely determined by $\{e^{\prime}_{\lambda}\}_{\lambda\in\Lambda}$. Let $\tilde{e}$ be an $H$-admissible frame field of $N^{*}$ defined on an open set $U$ of $M$. Then for $\lambda\in\Lambda$ with $U\cap U_{\lambda}\not=\emptyset$, there exists an $H$-valued function $A^{\prime}_{\lambda}$ on $U\cap U_{\lambda}$ satisfying $\tilde{e}I^{\prime}_{4n,\varepsilon}=e^{\prime}_{\lambda}I^{\prime}_{4n,\varepsilon}A^{\prime}_{\lambda}$. This relation is equivalent to $\tilde{e}^{\prime}I^{\prime}_{4n,\varepsilon}=e_{\lambda}I^{\prime}_{4n,\varepsilon}A_{\lambda}$ with $A_{\lambda}:=I_{2n,2n}A^{\prime}_{\lambda}I_{2n,2n}$. Since $A_{\lambda}$ is $H$-valued, $\tilde{e}^{\prime}$ is an $H$-admissible frame field of $N$. Hence we obtain (b) in Theorem 4.5. Suppose that $N$ is an $H$-nilpotent structure of $(E,h,\nabla)$. Then we can suppose that the connection form $\omega_{\lambda}$ of $\nabla$ with respect to $e_{\lambda}I^{\prime}_{4n,\varepsilon}$ is valued in the Lie algebra of $H$. Therefore the connection form $\omega^{\prime}_{\lambda}$ of $\nabla$ with respect to $e^{\prime}_{\lambda}I^{\prime}_{4n,\varepsilon}$ is valued in the Lie algebra of $H$. Hence we obtain (c) in Theorem 4.5. $\square$
Let $N$ be an $H$-nilpotent structure of $(E,h)$. Then we call $N^{*}$ as in Theorem 4.5 the dual $H$-nilpotent structure of $N$. We see that $N$ is the dual $H$-nilpotent structure of $N^{*}$: $(N^{*})^{*}=N$.
@@
$I:=\dfrac{1}{2}(N+N^{*}),\quad J_{2}:=\dfrac{1}{2}(N-N^{*}),\quad J_{1}:=-IJ_{2}.$ (4.5)
@@
Let $I$ be a complex structure of $E$ and $J_{1}$, $J_{2}$ paracomplex structures of $E$. Suppose that $I$, $J_{1}$, $J_{2}$ give a paraquaternionic structure $E_{I,J_{1},J_{2}}$ such that $h$ is adapted to $E_{I,J_{1},J_{2}}$. Then $N:=I+J_{2}$ is an $H$-nilpotent structure of $(E,h)$ and $N^{*}:=I-J_{2}$ is its dual $H$-nilpotent structure.

### v3

@@ 4 Special nilpotent structures
Suppose that $\nabla$ is flat. Let $K$ be a Lie subgroup of $G$. Then for a $K$-nilpotent structure $N$ of $(E,h,\nabla)$, there exists a $K$-admissible frame field $e$ of $N$ on a neighborhood of each point of $M$ such that each local section of $E$ which appears in $e$ is parallel with respect to $\nabla$.
@@
Let $W$ be a $2n$-dimensional subspace of $E^{4n}_{2n}$ spanned by $\eta_{1},\dots,\eta_{n},\eta_{n+1},\dots,\eta_{2n}$. Then $W$ is a light-like subspace of $E^{4n}_{2n}$. Let $SO(2n,2n)_{W}$ be a Lie subgroup of $SO(2n,2n)$ which consists of elements preserving $W$. Then we have $G\subset SO(2n,2n)_{W}$ and $G\not=SO(2n,2n)_{W}$ ([4]). An $SO(2n,2n)_{W}$-nilpotent structure of $(E,h,\nabla)$ is just an $\varepsilon$-nilpotent structure of $(E,h)$ satisfying the Walker condition. If $K=SO(2n,2n)_{W}$, then the conclusion of Proposition 4.2 is not necessarily valid, because of Example 3.1 $\sim$ Example 3.4, while for an almost $\varepsilon$-nilpotent structure $N$ of $E^{4n}_{2n}$ satisfying the Walker condition with respect to the Levi-Civita connection, $\pi_{N}$ gives a constant $2n$-dimensional light-like subspace of $E^{4n}_{2n}$.
@@
See [7], [21] for paraquaternionic structures (notice the sign in (i)). We say that $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$ if $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ satisfy (i), (ii) and
@@
    the section of $E_{I,J_{1},J_{2}}$ given by $N_{r,\theta}:=r(I-(\sin\theta)J_{1}+(\cos\theta)J_{2})$ for $r\in\mbox{\boldmath{$R$}}\setminus\{0\}$ and $\theta\in[0,2\pi)$ is an $H$-nilpotent structure of $(E,h);$
@@
    There exists a unique light-like subbundle $\pi^{\times}_{N}$ of $(E,h)$ of rank $2n$ which is locally spanned by
@@
    for any $H$-admissible frame field $e$ of $N$. In particular, $E=\pi_{N}\oplus\pi^{\times}_{N}$.
@@
    There exists a unique $\varepsilon$-nilpotent and $H$-nilpotent structure $N^{\times}$ of $(E,h)$ such that each $H$-admissible frame field of $N^{\times}$ is given by $e^{\prime}=(e_{1},\dots,e_{2n},-e_{2n+1},\dots,-e_{4n})$ for an $H$-admissible frame field $e$ of $N$. In particular, $\pi_{N^{\times}}=\pi^{\times}_{N}$.
@@
    The condition that $N$ is an $H$-nilpotent structure of $(E,h,\nabla)$ is equivalent to the condition that $N^{\times}$ is an $H$-nilpotent structure of $(E,h,\nabla)$.
Proof  For an $H$-admissible frame field $e$ of $N$, the subbundle of the restriction of $E$ on the domain of $e$ spanned by $\xi^{\prime}_{i}$, $\xi^{\prime}_{n+i}$ ($i=1,\dots,n$) as in (4.4) does not depend on the choice of $e$, because of the definition of $H$ in (4.1). Therefore there exists a unique light-like subbundle $\pi^{\times}_{N}$ of $(E,h)$ of rank $2n$ satisfying (a) in Theorem 4.5. Hence we obtain (a) in Theorem 4.5. For $N$, we can find an open cover $\{U_{\lambda}\}_{\lambda\in\Lambda}$ of $M$ and a family $\{e_{\lambda}\}_{\lambda\in\Lambda}$ of $H$-admissible frame fields of $N$ satisfying (i), (ii) in Definition 4.1 for $K=H$. We set
@@
for $\lambda$, $\mu\in\Lambda$ satisfying $U_{\lambda}\cap U_{\mu}\not=\emptyset$. Then $A^{\prime}_{\lambda\mu}$ satisfies $e^{\prime}_{\mu}I^{\prime}_{4n,\varepsilon}=e^{\prime}_{\lambda}I^{\prime}_{4n,\varepsilon}A^{\prime}_{\lambda\mu}$ on $U_{\lambda}\cap U_{\mu}$, and since $A_{\lambda\mu}$ is valued in $H$, $A^{\prime}_{\lambda\mu}$ is also valued in $H$. Therefore there exists an $\varepsilon$-nilpotent and $H$-nilpotent structure $N^{\times}$ of $(E,h)$ such that $e^{\prime}_{\lambda}$ ($\lambda\in\Lambda$) are $H$-admissible frame fields of $N^{\times}$ and then $N^{\times}$ is uniquely determined by $\{e^{\prime}_{\lambda}\}_{\lambda\in\Lambda}$. Let $\tilde{e}$ be an $H$-admissible frame field of $N^{\times}$ defined on an open set $U$ of $M$. Then for $\lambda\in\Lambda$ with $U\cap U_{\lambda}\not=\emptyset$, there exists an $H$-valued function $A^{\prime}_{\lambda}$ on $U\cap U_{\lambda}$ satisfying $\tilde{e}I^{\prime}_{4n,\varepsilon}=e^{\prime}_{\lambda}I^{\prime}_{4n,\varepsilon}A^{\prime}_{\lambda}$. This relation is equivalent to $\tilde{e}^{\prime}I^{\prime}_{4n,\varepsilon}=e_{\lambda}I^{\prime}_{4n,\varepsilon}A_{\lambda}$ with $A_{\lambda}:=I_{2n,2n}A^{\prime}_{\lambda}I_{2n,2n}$. Since $A_{\lambda}$ is $H$-valued, $\tilde{e}^{\prime}$ is an $H$-admissible frame field of $N$. Hence we obtain (b) in Theorem 4.5. Suppose that $N$ is an $H$-nilpotent structure of $(E,h,\nabla)$. Then we can suppose that the connection form $\omega_{\lambda}$ of $\nabla$ with respect to $e_{\lambda}I^{\prime}_{4n,\varepsilon}$ is valued in the Lie algebra of $H$. Therefore the connection form $\omega^{\prime}_{\lambda}$ of $\nabla$ with respect to $e^{\prime}_{\lambda}I^{\prime}_{4n,\varepsilon}$ is valued in the Lie algebra of $H$. This means that $N^{\times}$ is an $H$-nilpotent structure of $(E,h,\nabla)$. Similarly, we can prove the converse. Hence we obtain (c) in Theorem 4.5. $\square$
Let $N$ be an $H$-nilpotent structure of $(E,h)$. Then we call $N^{\times}$ as in Theorem 4.5 the dual $H$-nilpotent structure of $N$. We see that $N$ is the dual $H$-nilpotent structure of $N^{\times}$: $(N^{\times})^{\times}=N$.
@@
$I:=\dfrac{1}{2}(N+N^{\times}),\quad J_{2}:=\dfrac{1}{2}(N-N^{\times}),\quad J_{1}:=-IJ_{2}.$ (4.5)
@@
Let $E$, $h$ be as in the beginning of Section 2. Let $I$ be a section of ${\rm End}\,E$. We say that $I$ is an $\varepsilon$-complex structure of $(E,h)$ if $I$ satisfies
  • (i)
    $I$ is a complex structure of $E$,
  • (ii)
    $I$ is $h$-preserving, that is, $I^{*}h=h$,
  • (iii)
    on a neighborhood of each point of $M$, there exists an ordered pseudo-orthonormal local frame field $e=(e_{1},\dots e_{2n},e_{2n+1},\dots,e_{4n})$ of $E$ giving the orientation and satisfying
    $IeI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}\Lambda_{n,+},\quad\Lambda_{n,+}:=\left[\begin{array}[]{cccc}O_{n}&-I_{n}&O_{n}&O_{n}\\ I_{n}&O_{n}&O_{n}&O_{n}\\ O_{n}&O_{n}&O_{n}&-I_{n}\\ O_{n}&O_{n}&I_{n}&O_{n}\end{array}\right].$
Let $I$ be an $\varepsilon$-complex structure of $(E,h)$. Then such a frame field as $e$ is called an admissible frame field of $I$. We see that $I$ gives a section $I^{*}$ of ${\rm End}\,E^{*}$ by $I^{*}\phi^{*}=\phi^{*}\circ I$ for a local section $\phi^{*}$ of $E^{*}$. Then we have $I^{*}e^{*}I^{\prime}_{4n,\varepsilon}=e^{*}I^{\prime}_{4n,\varepsilon}{}^{t}\!\Lambda_{n,+}$. Therefore, if we set $\Theta_{I}(\phi^{*},\psi^{*}):=h^{*}(\phi^{*},I^{*}\psi^{*})$, then $\Theta_{I}$ is a section of $\bigwedge^{2}\!E$ and locally represented as
$\Theta_{I}=\sum^{n}_{i=1}(e_{i}\wedge e_{n+i}-(\varepsilon 1)^{\delta_{i1}}e_{2n+i}\wedge e_{3n+i}).$ (4.6)
For an $h$-connection $\nabla$ of $E$, $\nabla I=0$ is equivalent to ^$\hat{\nabla}\Theta_{I}=0$. Let $N$ be an $\varepsilon$-nilpotent and $H$-nilpotent structure of $(E,h)$ and $e$ an $H$-admissible frame field of $N$. Let $I$, $J_{2}$ be as in (4.5). Then $I$ is an $\varepsilon$-complex structure of $(E,h)$ such that $e$ is an admissible frame field of $I$ and $J_{2}$ is an $\varepsilon$-paracomplex structure of $(E,h)$ in the sense of Remark 3.4 such that
$(e_{n+1},\dots,e_{2n},-e_{1},\dots,-e_{n},e_{2n+1},\dots,e_{3n},e_{3n+1},\dots,e_{4n})$
is an admissible frame field of $J_{2}$. In addition, $\Theta_{N}=\Theta_{I}+\Theta_{J_{2}}$. Let $J_{1}$ be as in (4.5). Then $J_{1}$ is an $\varepsilon$-paracomplex structure of $(E,h)$ in the sense of Remark 3.4 such that $e$ is an admissible frame field of $J_{1}$.
Remark 4.4
Let $I$ be a complex structure of $E$ and $J_{1}$, $J_{2}$ paracomplex structures of $E$. Suppose that $I$, $J_{1}$, $J_{2}$ give a paraquaternionic structure $E_{I,J_{1},J_{2}}$ such that $h$ is adapted to $E_{I,J_{1},J_{2}}$. Then $N:=I+J_{2}$ is an $H$-nilpotent structure of $(E,h)$ and $N^{\times}:=I-J_{2}$ is its dual $H$-nilpotent structure.